What is the present value of the following? Use Appendix B as an approximate answer, but calculate your final answer using the formula and financial calculator methods. a. $7,600 in 12 years at 5 percent? (Do not round intermediate calculations. Round your final answer to 2 decimal places.) b. $16,300 in 6 years at 8 percent? (Do not round intermediate calculations. Round your final answer to 2 decimal places.) c. $25,500 in 18 years at 7 percent? (Do not round intermediate calculations. Round your final answer to 2 decimal places.) HintseBook & Resources Hint #1
Answer #1

Present value=Cash flow*Present value of discounting factor(rate%,time period)

a.Present value=7600/(1.05)^12

=$4231.96(Approx).

b.Present value=16300/(1.08)^6

=$10271.76(Approx).

c.Present value=25500/(1.07)^18

=$7544.53(Approx).